OpenStudy (anonymous):

Replace the variables a and b with integers and solve for log a b = x by using the change of base formula. Show all work and use complete sentences to explain your steps in solving the logarithm you created.

OpenStudy (zepp):

Let's say a = a and b = b

OpenStudy (zepp):

And \(\large \log_ab=x\) Rewrite is in exponential form. \(\large a^x = b\) Take the log of the both side \(\large log_ax= \log b\) By using a log property \(\large x log_a = log_b\) Divide by log a to both side \(\large x = \frac{\log b}{\log a}\)

OpenStudy (dumbcow):

\[\log_{a} b = \frac{\log b}{\log a}\]

OpenStudy (zepp):

Therefore, the equation @dumbcow typed.

OpenStudy (anonymous):

but dont i need numbers instead of variables?

OpenStudy (dumbcow):

yes , @zepp showed the derivation of where the change of base formula came from

OpenStudy (zepp):

The property may not be very clear but there's a hidden base.

OpenStudy (zepp):

That's why I can apply the property.

OpenStudy (zepp):

Arrgh, I lost everything I just typed :(

OpenStudy (zepp):

\(\huge \begin {align} a^x&=b \\ log_{10}a^x &= log_{10}b \\ xlog_{10}a&=log_{10}b \\ x &= \frac{\log_{10}b}{log_{10}a} \end {align}\) There, we can ignore the base, but that's just to show the invisible base there.

OpenStudy (anonymous):

thanks